Find the general solution of the Lagrange's equation 2yzp + zxq = 3xy

1
Expert's answer
2022-01-12T14:37:27-0500

"\frac{dx}{2yz}=\frac{dy}{zx}=\frac{dz}{3xy}"


"3ydy=zdz"

"3y^2-z^2=c_1"


"xdx=2ydy"

"x^2-2y^2=c_2"


"F(c_1,c_2)=F(3y^2-z^2,x^2-2y^2)=0"